OpenStudy (anonymous):

Which of these equations represent functions? A. y= x^2 B. y^2= x^2 C. x^2-y^2=1 D. y= 3x

OpenStudy (anonymous):

Ah, well let us begin by defining what is a function and what is not a function, shall we?

OpenStudy (anonymous):

A function is a mapping from a set of values to another set of values so that every input has exactly one output. Two inputs may be mapped to the same output, but in no case is it allowed that 1 input be mapped to two separate outputs.

OpenStudy (anonymous):

Following me?

OpenStudy (anonymous):

I am now

OpenStudy (anonymous):

Okay good.

OpenStudy (anonymous):

So let's look at the functions given. The first one is y = x^2 Does it meet the requirements of a function? Is there any x that gets mapped to 2 different ys?

OpenStudy (anonymous):

I don't think there is

OpenStudy (anonymous):

If the x is mapped to two different ys, it's not a function, right? Or is it the other way around?

OpenStudy (anonymous):

Correct. If even one x gets mapped to two different ys, then it's not a function.

OpenStudy (anonymous):

And like you said, for that first part, there are no xs that get mapped to two ys.

OpenStudy (anonymous):

So would A be one of the correct answers?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

So now let's look at B. y^2 = x^2.

OpenStudy (anonymous):

How can I determine by looking at the equation whether or not one x maps to two ys?

OpenStudy (anonymous):

Well... just start by looking at an x value or two. Pick x = 2 for example.

OpenStudy (anonymous):

x^2 = 4, so are there multiple ys or just 1 y that satisfy the equation y^2 = 4?

OpenStudy (anonymous):

There's only one y in the equation for B, so that works as a function, right?

OpenStudy (anonymous):

I'm thinking that all except D are correct

OpenStudy (anonymous):

Let's slow down and look at B.

OpenStudy (anonymous):

So for x = 2, y = 2 satisfies the equation y^2 = x^2, right?

OpenStudy (anonymous):

I'm thinking yes

OpenStudy (anonymous):

Yes. It does. So 2 gets mapped to 2.

OpenStudy (anonymous):

But is there another y so that y^2 = 2^2?

OpenStudy (anonymous):

I don't see another y

OpenStudy (anonymous):

So it only gets mapped to one y

OpenStudy (anonymous):

No? Tell me. What if y = -2? What is y^2 then?

OpenStudy (anonymous):

-4?

OpenStudy (anonymous):

I'm not very good at Algebra 2, lol

OpenStudy (anonymous):

(-2)*(-2) = -4? You're multiplying two negative numbers.

OpenStudy (anonymous):

You're doing great. Just stick with me =) Don't get discouraged.

OpenStudy (anonymous):

So it would be a positive 4 instead?

OpenStudy (anonymous):

Yes. So for y= -2, y^2 = 4 = x^2, so y^2 = x^2, right?

OpenStudy (anonymous):

Which means what?

OpenStudy (anonymous):

There's more than one y, so it wouldn't be a function?

OpenStudy (anonymous):

Yes. So we looked at x = 2, saw that x^2 =4 and wondered what y values would work then. And we saw that for both y = 2 and y = -2, y^2 = x^2 worked. Which means that x = 2 gets mapped to both 2 and -2.

OpenStudy (anonymous):

Get it?

OpenStudy (anonymous):

I think so. So B isn't an answer

OpenStudy (anonymous):

Right. Let's look at C.

OpenStudy (anonymous):

x^2 - y^2 = 1

OpenStudy (anonymous):

Look at it for a bit, maybe at one or two xs and tell me what you think.

OpenStudy (anonymous):

I'm not coming up with anything by looking at it, lol.

OpenStudy (anonymous):

Wait, there would be 2 y's since the y is squared?

OpenStudy (anonymous):

Yes. Let's see if we can come up with an example.

OpenStudy (anonymous):

What about when x is 2? x^2 - y^2 = 1 so 4 - y^2 = 1 so -y^2 = -3 so y^2 = 3 What values for y work in this case?

OpenStudy (anonymous):

sqrt(3) works. And?